on a final examination in biology, the mean was 75 and the standard deviation was 12. determine the standard score of a student who received a score of 60 assuming that the scores are normally distributed.


1
Expert's answer
2022-03-01T17:53:48-0500

Solution:

"z=\dfrac{x-\mu}{\sigma}=\dfrac{60-75}{12}=-1.25"